LASER-INDUCED-FLUORESCENCE EXCITATION AND EMISSION SPECTRA OF THE 1^{3}Σ_{u}^{+}–1^{3}Π_{g} (1^{3}Σ^{+}–2^{3}Π) TRANSITION IN K_{2}, Rb_{2} (KRb) ON THE SURFACE OF SUPERFLUID HELIUM DROPLETS

\underline{OLIVIER ALLARD}, GERALD AUBÖCK, JOHANN NAGL, CARLO CALLEGARI and WOLF-GANG E. ERNST, Institute of Experimental Physics, TU Graz, Petersgasse 16, 8010 Graz, Austria/EU.

We measured laser induced fluorescence and emission spectra of the 1^{3}Σ_{u}^{+}(0_{u}^{−}, 1_{u})–1^{3}Π_{g}(0_{g}^{+}, 0_{g}^{−}, 1_{g}, 2_{g}) transition of K_{2}, Rb_{2}, and the corresponding 1^{3}Σ^{+}(0^{−}, 1)–2^{3}Π(0^{+}, 0^{−}, 1, 2) transition of KRb, on the surface of superfluid helium droplets. The spectra are relatively well explained, quantitatively, by simulations making use of the gas phase potentials.^{a}^{b} For an exact agreement, however, one needs to assume: (1) that the 1 ← 0^{−} excitation is strongly suppressed, (2) that the spin-orbit coupling constant is increased (K_{2}, KRb), and (3) an adjustable broadening factor. (1) points to a yetunexplained lack of population of the 0^{−} state after dimer formation, (2) and (3) can be justified qualitatively as due to the presence of the droplet, but are difficult to predict quantitatively. Data analysis is in progress. Emission spectra, sharp, confirm that the molecules emit from the gas phase after having been ejected from the droplet; they can be used to test existing potential energy surfaces. We are in the process of acquiring new data and improving our analysis of vibrational populations in the excited state.^{c}

aK_{2}: A. R. Allouche. Potential energy curves available at lasim.univ-lyon1.fr/allouche/pec.html

bRb_{2}, KRb: M. Aymar and O. Dulieu private communication

cJ. Nagl, G. Auböck, C. Callegari, and W. E. Ernst, paper TC05, 60^{th} OSU Symposium on Molecular Spectroscopy, 2005